OpenStudy (anonymous):

Find the height of a cylinder with radius 5 cm and surface area 1000 square cm.

hartnn (hartnn):

Imagine a cylinder  it has 2 bases and one curved surface. Base area = \(\pi r^2 \) curved area = \(h(2\pi r)\) so total surface area = \(2\pi r^2 +2\pi rh \)